what is the mass of the precipitate formed when 50ml of 16.9%solution of AgNO3 is mixed with 50ml of 5.8%NaCl solution

Grade:12th pass

2 Answers

Vikas TU
4 years ago
Hiii 
W/V = Mass of solute / Vol of Solution 
Ag +Cl = Agcl
Massof silver nitrate = 16.9 * 50 ml = 8.45 g
Molar quantity = 8.45 / 169.87 = 0.0497
Mass of Nacl = 2.90 g 
Molar quantity = 0.0497
AgNo3 + Nacl = Agcl + NaNo3
0.04974* 143.32 = 7.11g 
Arun
4 years ago
Moles of AgNo3= 50x 16.9/100x169.8
                           = 0.05mole
Moles of NaCl = 50x5.8/100x58.5
                       =0.05mole
AgNo3 + NaCl ----> AgCl + NaNo3
mass of AgCl = mole x molar mass
                       =0.05 x 143.5
                       = 7.16 g